# Scalable multi-qubit intrinsic gates in quantum dot arrays

### v1

@@ IV.1 The problem of bond inhomogeneity
Examining the bond strength condition, for example, of the CZZ gate,
@@
One way to think of the of intrinsic multi-control gates is to examine the “$k$-space” picture. Time evolution trace out a parametric path, which is a straight line with tangent vector $\bm{v}=\tfrac{1}{2\pi}(T_{1}-S_{1},T_{2}-S_{2})$ here. The target control gate is specified by a set of lattice points in the $k$-space, separating by distance 1. The goal for the time evolution is to bump into one of the lattice points, as illustrated in Fig. 5.
@@
Figure 5: (a) DC evolution path in the (extended) $k$-space. The red lattice points represent the target control gate and straight lines represent different evolution paths: the ideal case with identical bonds (thick green line), bonds with mutually rational strengths (dashed green line) and bonds with mutually irrational strengths (thin blue line). (b) and (c), the reduced $k$-space time evolution for the rational and irrational bonds.
In the ideal case, the strength of edge-1 equals that of edge-2, $T_{1}-S_{1}=T_{2}-S_{2}$. In such case, the path would directly heads towards the first lattice point at $(1/2,1/2)$, as indicated by the thick green line in Fig. 5(a). This gives the theoretically shortest evolution time of
$\tau_{\mathrm{min}}=\frac{\pi}{T-S}$ (57)
this evolution time is the same as that of two-qubit CZ gate, despite more target qubits are involved.
In reality, due to structural defects and imperfections, it is hard to have completely identical bond strength. Although it is possible to calibrate the bond strength for small degree of freedoms, the issue of inhomogeneity poses a major challenge when the size of qubit chips scale up. From our theory, different coupling strengths are responsible to the different velocity complement of the $k$-space trajectory. And the time evolution target can be any lattice points offsets by a set of integers. Therefore, even if the bonds are inhomogeneous, it is still possible that the evolution path would bump into some lattice point other than $(1/2,1/2)$. One such example is plotted with dashed green line in Fig. 5(a), where the time evolution reaches the lattice point $(\frac{5}{2},\frac{3}{2})$. Apparently, the condition for perfectly attaining a lattice points is that the effective strength $T-S$ between different bonds being rational to each other. In general, the bond strength ratio is irrational, and the evolution path will never perfectly arrive at a lattice point, as illustrated with the thin blue line in Fig. 5(a). This imperfection leads to an error in implementing the target gate. One can still argue that given sufficiently long evolution time, the evolution path can get arbitrarily close to the target point. To understand this property, we consider the reduced $k$-space indicated by the shaded area in Fig. 5(a). Due to the periodic nature of the target gate, we can fold a full evolution path into the first lattice by simply discarding the integer parts of its vector components. This process is illustrated in Fig. 5(b) and Fig. 5(c) for the rational green path and irrational blue path. For the rational path, the trajectory attain the target point after finite cycles, while for the irrational path, the trajectory will densely cover the whole reduced $k$-space given enough time.
Although the above discussion is made for a three-qubit system, it holds true for general $n$-qubit multi control gates. One simply need to replace the two-dimensional $k$-space picture with a ${n-1}$-dimensional space. We can proceed to estimate an upper bond of the evolution time given target gate infidelity $\epsilon$ by
$\tau\simeq\frac{1}{(n-1)^{1/2}v_{m1}}\Gamma(\frac{n}{2})\left(\frac{8/\pi}{\epsilon(n-1)}\right)^{(n-2)/2}.$ (58)
The result is derived by transforming the evolution path into a tube with finite cross-section and ask when does the tube fill the entire reduced $k$-space [63].
Despite that eventually the target can be achieved. The problem is that the time would take too long following a simple straight path. It is very likely the during these time, the qubit will already decohere. Therefore, one must find a way to dramatically accelerate the time evolution. It turns out one can design a controlled path pof evolution by pulsing the target qubits.
IV.2 Dynamical calibration: example
The time requirement scales exponentially when the bonds are inhomogeneous. We propose a scheme that will effectively eliminate this issue by “steering the path” in the phase space with fast single-qubit pulses.
@@ IV.2 Dynamical calibration: example
Till now, we have only considered time evolutions with time-independent control, i.e., the exchange coupling is turned on and stays constant during the full evolution. This is often termed as DC gates in literatures. We denote the rotating-frame quantum gate at time $\tau$ as
@@
Going beyond the simple DC-control scheme, let us consider applying a bit-flip pulse $X_{j}$ to the $j$-th qubit at time $\tau_{1}<\tau$. Assuming the pulse width to be much shorter compared to the time duration for achieving the full multi-qubit gate, we can apply a similar perturbative treatment to calculate the resulting quantum gate
@@
where $|\xi_{w}^{\prime}\rangle\equiv X_{j}|\xi_{w}\rangle$ is the flipped entangled state containing dot $j$. Effectively, the $X_{j}$ pulse swap the spin-conserved and spin-flipped tunneling strength $T_{w}=J\lvert t_{w}\rvert^{2}$ and $S_{w}=J_{w}\lvert s_{w}\rvert^{2}$. For multi-control gates, exchanging $T$ and $S$ tunneling channel of a bond results in a reversal of its effective velocity component. This offers a control possibility where we can design an customize path in the $k$-space to the target gate by applying single qubit pulses at appropriate intervals.
@@
$|\xi_{\mathsf{c}j}\rangle=s^{*}|\uparrow\uparrow\rangle_{\mathsf{c}j}+t^{*}|\uparrow\downarrow\rangle_{\mathsf{c}j}-t\,|\downarrow\uparrow\rangle_{\mathsf{c}j}+s\,|\downarrow\downarrow\rangle_{\mathsf{c}j},$ (63)
where we omit the normalization factor ${1}/{\sqrt{2}}$ and the subscripts for $t$ and $s$. Applying bitflip to the target or control qubit results in,
$X_{j}|\xi_{\mathsf{c}j}\rangle$ $=t^{*}|\uparrow\uparrow\rangle_{\mathsf{c}j}+s^{*}|\uparrow\downarrow\rangle_{\mathsf{c}j}+s\,|\downarrow\uparrow\rangle_{\mathsf{c}j}-t\,|\downarrow\downarrow\rangle_{\mathsf{c}j},$ (64)
$X_{\mathsf{c}}|\xi_{\mathsf{c}j}\rangle$ $=-t\,|\uparrow\uparrow\rangle_{\mathsf{c}j}\!+s\,|\uparrow\downarrow\rangle_{\mathsf{c}j}+s^{*}|\downarrow\uparrow\rangle_{\mathsf{c}j}+t^{*}|\downarrow\downarrow\rangle_{\mathsf{c}j}.$
In both cases, the spin-conserved and spin-flipped tunneling strength $T$ and $S$ are exchanged. In terms of the $k$-space velocity $\bm{v}=(v_{1},v_{2})$ in Eq. (56), the $X_{j}$ pulse flips the velocity component $v_{j}\to-v_{j}$ while the $X_{\mathsf{c}}$ pulse flips both velocity components. Supposing $v_{1}>v_{2}>0$, the simplest scheme to overcome bond inhomogeneity is to apply a $X_{1}$ pulse at time instant $\tau_{1}=\frac{1}{4}(v_{1}^{-1}+v_{2}^{-1})$, as illustrated by the path in Fig. 6(a). The total gate time $\tau={1}/{2v_{2}}$ is determined by the weaker bond of the two. Alternatively, one can apply two $X_{1}$ pulses symmetrically to cancel out the effect of the remaining $X$ gate, show by the path in Fig. 6(b). One step further, one can apply pulses to all three qubits in the order of $X_{1}$, $X_{2}$ and $X_{\mathsf{c}}$, separated by time duration $\tau_{1}$, $\tau_{2}$ and $\tau_{3}$ respectively. The corresponding rotating-frame evolution can be calculated to be
@@
$\quad\simeq X_{\mathsf{c}}X_{2}X_{1}\,\mathrm{e}^{-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\phi^{\prime}_{\mathsf{c}}\sigma_{\mathsf{c}}^{Z}}\mathrm{e}^{-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\phi^{\prime}_{2}\sigma_{2}^{Z}}\mathrm{e}^{-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\phi^{\prime}_{1}\sigma_{1}^{Z}}\mathrm{e}^{\mkern 1.0mu\mathrm{i}\mkern 1.0mu\sum_{k=1}^{4}\tau_{k}\bm{\Lambda}_{k}},$ (65)
where apart from the single-qubit bit-flips $X_{\mathsf{c}}X_{2}X_{1}$, the pulses induce extra local phases $\phi_{\mathsf{c}}^{\prime}=(\tau_{1}+\tau_{2}+\tau_{3})\varepsilon_{\mathsf{c}}^{\mathrm{Z}}$, $\phi_{2}^{\prime}=(\tau_{1}+\tau_{2})\varepsilon_{2}^{\mathrm{Z}}$ and $\phi_{1}^{\prime}=\tau_{1}\varepsilon_{1}^{\mathrm{Z}}$. The entangling part is determined by a four-step time evolution, under the effective grid vectors $\bm{\Lambda}_{1}=\bm{\Lambda}$, $\bm{\Lambda}_{2}=X_{1}\bm{\Lambda}X_{1}$, $\bm{\Lambda}_{3}=X_{2}X_{1}\bm{\Lambda}X_{1}X_{2}$, and $\bm{\Lambda}_{4}=X_{\mathsf{c}}X_{2}X_{1}\bm{\Lambda}X_{1}X_{2}X_{\mathsf{c}}$ for the time duration of $\tau_{1},\tau_{2},\tau_{3}$, and $\tau_{4}\equiv\tau-(\tau_{1}+\tau_{2}+\tau_{2})$ respectively. The $k$-space path is illustrated in Fig. 6(c).
Compared with a single $X_{1}$ pulse, one key advantage of pulsing all three qubits is its potential to be combined with dynamical decoupling (DD), where all three qubits can be protected by decoupling pulses while performing the gate. Specially, let us consider applying the $XYXY$ sequence to each of the qubits. For the system-environment coupling Hamiltonian of $H_{\mathrm{SB}}$, the net effect of the sequence on a single qubit is
@@
The Pauli group twirling results in identity on the system part, hence the pulse sequence can suppress noise effect up to the first order in system-environment coupling. Similar to the $X$-pulse, a $Y$-pulse acts on the entangled state by
$Y_{j}|\xi_{\mathsf{c}j}\rangle$ $=\mkern 1.0mu\mathrm{i}\mkern 1.0mut|\uparrow\uparrow\rangle_{\mathsf{c}j}-\mkern 1.0mu\mathrm{i}\mkern 1.0mus|\uparrow\downarrow\rangle_{\mathsf{c}j}+\mkern 1.0mu\mathrm{i}\mkern 1.0mus^{*}\,|\downarrow\uparrow\rangle_{\mathsf{c}j}+\mkern 1.0mu\mathrm{i}\mkern 1.0mut^{*}\,|\downarrow\downarrow\rangle_{\mathsf{c}j},$ (67)
$Y_{\mathsf{c}}|\xi_{\mathsf{c}j}\rangle$ $=-\mkern 1.0mu\mathrm{i}\mkern 1.0mut^{*}|\uparrow\uparrow\rangle_{\mathsf{c}j}+\mkern 1.0mu\mathrm{i}\mkern 1.0mus^{*}|\uparrow\downarrow\rangle_{\mathsf{c}j}-\mkern 1.0mu\mathrm{i}\mkern 1.0mus|\downarrow\uparrow\rangle_{\mathsf{c}j}-\mkern 1.0mu\mathrm{i}\mkern 1.0mut|\downarrow\downarrow\rangle_{\mathsf{c}j}.$
which also flips the $T$ and $S$ tunneling strength of a particular bond. Based on this similarity between $X$ and $Y$ pulses, we can devise a dynamical calibration scheme that involves interleaved $X$ and $Y$ pulses. The pulsing choice and timing breakdown is illustrated in Fig. 6(e). This dynamical calibration seems like a trivial repetition of the four-staged evolution with $X_{1}$,$X_{2}$ and $X_{3}$ pulses in the $k$-space, see Fig. 6(d). However, the reduced pulse sequence for each qubit is universal dynamical decoupling sequence. Therefore, each qubit is decoupled from the environmental interaction by the pulses, while these qubits remain coherently coupled to each other so that a MQCP gate can be performed on them. Hence we see that our dynamical calibration scheme not only is capable of mitigating the inhomogeneity errors, but also allow enhancing the overall qubit quality.

### v2

@@ IV.1 The problem of bond inhomogeneity
For MQCP gates, the dynamics rule Eq. (38) specifies the condition that must be met by the connecting bonds in order to implement the target gate. Examining the dynamics rule, for example, of the $\mathrm{CZ}_{1}Z_{2}$ gate,
@@
One way to think of the gate dynamics is to examine the “phase space” picture. As as illustrated in Fig. 5, the target gate is specified by a lattice of half-integer points. Time evolution traces out a parametric path of $\tau$, which is a straight line with the tangent vector $\bm{v}\propto(T_{1}-S_{1},T_{2}-S_{2})$. The goal for time evolution is to bump into one of the lattice points. In the ideal case, $T_{1}-S_{1}=T_{2}-S_{2}$ and the path would directly heads towards the first lattice point at $(1/2,1/2)$. This corresponds to the theoretically minimal evolution time of $\tau={\pi}/{(T-S)}$, which is the same as that of the CZ gate in despite of more target qubits. In reality, due to structural defects and imperfections, it is unlikely to have completely identical interdot couplings. The difference in the two bonds results in unwanted error in implementing the target gate.
As the target can be any lattice point, it is still possible for the evolution path to bump into some lattice point even if the bonds are inhomogeneous. However, for perfectly attaining a lattice point, the ratio for velocity components must be rational. For irrational velocity ratios, the evolution path can never come across a lattice point, as illustrated with the thin blue line in Fig. 5(a). Instead, the evolution path can get arbitrarily close to a lattice point given sufficiently long time. To demonstrate this property, we define the “reduced” phase space as the first lattice indicated by the shaded area in Fig. 5(a). Due to the periodic nature of the target gate, we can fold the full evolution path into the reduced phase space. The results are illustrated in Fig. 5(b) and Fig. 5(c) for the rational and irrational path. For the rational path, the trajectory attains the target point after finite cycles. For the irrational path, the trajectory will densely covers the reduced phase space given sufficient time.
@@
Figure 5: (a) DC evolution path in the (extended) phase space. The red lattice points represent the target gate and straight lines represent different evolution paths: the ideal case with identical bonds (thick green line), bonds with mutually rational strengths (dashed green line) and bonds with mutually irrational strengths (thin blue line). (b) and (c), the reduced phase space time evolution for the rational and irrational bonds.
Although the above discussion is based on the $\mathrm{CZ}_{1}Z_{2}$ gate, similar conclusions hold for a general $(N+1)$-qubit MQCP gate. One only need to replace the two-dimensional picture with a similar ${N}$-dimensional version, and the “velocity” vector by $\bm{v}=(T_{1}-S_{1},T_{2}-S_{2},\cdots,T_{N}-S_{N})$, with different coupling strengths corresponding to different velocity complements. Despite that the target point can be eventually approached, the time would take too long for large systems. In fact, the worst-case time required to achieve gate infidelity $\epsilon$ scales super-exponentially with $N$,
$\tau\propto\Gamma(\frac{N}{2})\Bigl{(}\frac{8}{\epsilon\pi N}\Bigr{)}^{(N-1)/2},$ (62)
where $\Gamma(\cdot)$ is the gamma function [68]. It is very likely the during these time, the qubit will decohere beyond recognition! Although it is possible to fine-tune the bond strength for small systems, the issue of inhomogeneity poses a major challenge when the size of the qubit chip scales up. For scalable multi-qubit gates, one must find an efficient way to overcome the bond inhomogeneity.
We propose a control scheme that can effectively eliminate the bond inhomogeneity issue. The grand idea is to “steer an artificial path” in the phase space with external pulses. Inspired by the dynamical decoupling protocol [52], we refer such protocol as dynamical calibration, in the sense that the bond strength are calibrated dynamically with pulses. The scheme scales well with the number of qubits, as only fast single-qubit pulses are required. In the followings, we will first explore this ideal with some examples, and then present a general theory of dynamical calibration.
IV.2 Dynamical calibration: examples
@@
Till now, we have only considered time evolutions with time-independent control, i.e., the exchange coupling is turned on and stays constant during the full evolution. This is often termed as DC gates [27]. So-far we have approximated the rotating-frame map at time $\tau$ with
@@
Going beyond the simple DC-control scheme, let us consider applying a bit-flip pulse $X_{j}$ to the $j$-th qubit at time $\tau_{1}<\tau$. Assuming the pulse width to be much shorter than the time duration for achieving the full multi-qubit gate, we can apply a similar perturbative treatment to calculate the resulting quantum gate
@@
where $|\xi_{w}^{\prime}\rangle\equiv X_{j}|\xi_{w}\rangle$ is the flipped entangled state containing dot $j$. Effectively, the $X_{j}$ pulse swap the spin-conserved and spin-flipped tunneling strength $T_{w}=J\lvert t_{w}\rvert^{2}$ and $S_{w}=J_{w}\lvert s_{w}\rvert^{2}$. For multi-control gates, exchanging $T$ and $S$ tunneling channel of a bond results in a reversal of its effective velocity component. This offers a control possibility where we can design an customize path in the phase space to the target gate by applying single qubit pulses at appropriate intervals.
@@
$|\xi_{\mathsf{c}j}\rangle=\tilde{s}\,|\uparrow\uparrow\rangle_{\mathsf{c}j}-\tilde{t}\,|\uparrow\downarrow\rangle_{\mathsf{c}j}+\tilde{t}^{*}\,|\downarrow\uparrow\rangle_{\mathsf{c}j}+\tilde{s}^{*}\,|\downarrow\downarrow\rangle_{\mathsf{c}j},$ (67)
where we omit the normalization factor ${1}/{\sqrt{2}}$ and the subscripts for $t$ and $s$ for brevity. Applying bitflip to the target or control qubit results in,
$X_{j}|\xi_{\mathsf{c}j}\rangle$ $=-\tilde{t}\,|\uparrow\uparrow\rangle_{\mathsf{c}j}+\tilde{s}\,|\uparrow\downarrow\rangle_{\mathsf{c}j}+\tilde{s}^{*}|\downarrow\uparrow\rangle_{\mathsf{c}j}+\tilde{t}\,|\downarrow\downarrow\rangle_{\mathsf{c}j},$ (68)
$X_{\mathsf{c}}|\xi_{\mathsf{c}j}\rangle$ $=\tilde{t}^{*}|\uparrow\uparrow\rangle_{\mathsf{c}j}+\tilde{s}^{*}|\uparrow\downarrow\rangle_{\mathsf{c}j}+\tilde{s}\,|\downarrow\uparrow\rangle_{\mathsf{c}j}-\tilde{t}\,|\downarrow\downarrow\rangle_{\mathsf{c}j}.$
In both cases, the relevant spin-conserved and spin-flipped tunneling strengths are exchanged $T\leftrightarrow S$. In terms of the phase space velocity $\bm{v}=(T_{1}-S_{1},T_{2}-S_{2})$, the $X_{j}$ pulse flips the velocity component $v_{j}\to-v_{j}$ while the $X_{\mathsf{c}}$ pulse flips both velocity components. Supposing $v_{1}>v_{2}>0$, the simplest scheme to overcome bond inhomogeneity is to apply a $X_{1}$ pulse at time instant $\tau_{1}=\frac{1}{4}(v_{1}^{-1}+v_{2}^{-1})$, as illustrated by the path in Fig. 6(a). The total gate time $\tau={1}/{2v_{2}}$ is determined by the weaker bond of the two. Alternatively, one can apply two $X_{1}$ pulses symmetrically to cancel out the effect of the remaining $X$ gate, show by the path in Fig. 6(b). One step further, one can apply pulses to all three qubits in the order of $X_{1}$, $X_{2}$ and $X_{\mathsf{c}}$, separated by time duration $\tau_{1}$, $\tau_{2}$ and $\tau_{3}$ respectively. The corresponding rotating-frame evolution can be calculated to be
@@
$\quad\simeq X_{\mathsf{c}}X_{2}X_{1}\,\mathrm{e}^{-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\phi^{\prime}_{\mathsf{c}}\sigma_{\mathsf{c}}^{Z}}\mathrm{e}^{-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\phi^{\prime}_{2}\sigma_{2}^{Z}}\mathrm{e}^{-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\phi^{\prime}_{1}\sigma_{1}^{Z}}\mathrm{e}^{\mkern 1.0mu\mathrm{i}\mkern 1.0mu\sum_{k=1}^{4}\tau_{k}\bm{\Lambda}_{k}}.$ (69)
Apart from single-qubit bit-flips $X_{\mathsf{c}}X_{2}X_{1}$, the pulses also induce local phase shifts $\phi_{\mathsf{c}}^{\prime}=(\tau_{1}+\tau_{2}+\tau_{3})\varepsilon_{\mathsf{c}}^{\mathrm{Z}}$, $\phi_{2}^{\prime}=(\tau_{1}+\tau_{2})\varepsilon_{2}^{\mathrm{Z}}$ and $\phi_{1}^{\prime}=\tau_{1}\varepsilon_{1}^{\mathrm{Z}}$, which are correctable by virtual-$Z$ gates. The entangling part is determined by a four-step time evolution, under the effective grid vectors $\bm{\Lambda}_{1}=\bm{\Lambda}$, $\bm{\Lambda}_{2}=X_{1}\bm{\Lambda}X_{1}$, $\bm{\Lambda}_{3}=X_{2}X_{1}\bm{\Lambda}X_{1}X_{2}$, and $\bm{\Lambda}_{4}=X_{\mathsf{c}}X_{2}X_{1}\bm{\Lambda}X_{1}X_{2}X_{\mathsf{c}}$ for the time duration of $\tau_{1},\tau_{2},\tau_{3}$, and $\tau_{4}\equiv\tau-(\tau_{1}+\tau_{2}+\tau_{2})$ respectively. The phase space path is illustrated in Fig. 6(c).
Compared with a single $X_{1}$ pulse, one key advantage of pulsing all three qubits is its potential to be combined with dynamical decoupling, where all three qubits can be protected by decoupling pulses while performing the gate. Specifically, let us consider applying the $XYXY$ sequence to each of the qubits. For the system-environment coupling Hamiltonian of $H_{\mathrm{SB}}$, the net effect of the sequence on a single qubit is
@@
The Pauli group twirling results in identity on the system part, hence the pulse sequence can suppress environmental coupling to the first order. Similar to the $X$-pulse, a $Y$-pulse acts on the entangled state by
$Y_{j}\,|\xi_{\mathsf{c}j}\rangle$ $=-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\tilde{t}^{*}|\uparrow\uparrow\rangle_{\mathsf{c}j}+\mkern 1.0mu\mathrm{i}\mkern 1.0mu\tilde{s}^{*}|\uparrow\downarrow\rangle_{\mathsf{c}j}-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\tilde{s}\,|\downarrow\uparrow\rangle_{\mathsf{c}j}-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\tilde{t}\,|\downarrow\downarrow\rangle_{\mathsf{c}j},$ (71)
$Y_{\mathsf{c}}\,|\xi_{\mathsf{c}j}\rangle$ $=\mkern 1.0mu\mathrm{i}\mkern 1.0mu\tilde{t}\,|\uparrow\uparrow\rangle_{\mathsf{c}j}-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\tilde{s}\,|\uparrow\downarrow\rangle_{\mathsf{c}j}+\mkern 1.0mu\mathrm{i}\mkern 1.0mu\tilde{s}^{*}|\downarrow\uparrow\rangle_{\mathsf{c}j}+\mkern 1.0mu\mathrm{i}\mkern 1.0mu\tilde{t}^{*}|\downarrow\downarrow\rangle_{\mathsf{c}j}.$
which also exchanges the $T$ and $S$ tunneling strengths of a particular bond. Based on this similarity between $X$ and $Y$ pulses, we can devise a dynamical calibration scheme that involves interleaved $X$ and $Y$ pulses. The pulsing choice and timing breakdown is illustrated in Fig. 6(e). This dynamical calibration seems like a trivial repetition of the four-staged evolution with $X_{1}$,$X_{2}$ and $X_{3}$ pulses in the phase space, see Fig. 6(d), but the reduced pulse sequence for each qubit is now a dynamical decoupling sequence. While the qubits are decoupled from the environment, they still remain coherently coupled to each other so that a MQCP gate can be performed. Hence we see that our dynamical calibration scheme not only is capable of mitigating the inhomogeneity error, but also enhances the overall qubit quality.